OpenStudy (anonymous):

the 6th term of an arithmetic sequence is 24. what is the sum of the 5th and 7th terms

OpenStudy (amistre64):

no way to determine without knowing the value of another term

OpenStudy (anonymous):

its 48

OpenStudy (amistre64):

24 = a1 + d(23) we have one equation in 2 unknowns

OpenStudy (anonymous):

let first term be a. let difference be d so for sixth term we have a+5d=24 so 5th term= a+4d 7th term= a+6d so 5th term +7th term = a+4d+a+6d =2a+10d =2(a+5d)=2*24 =48

OpenStudy (anonymous):

thx

OpenStudy (amistre64):

hmm, 48 does make sense ....

OpenStudy (anonymous):

in ap difference is same between consecutive nos so, T6-T5=T7-T6 2T6=T5+T7 2*24=48 48 IS ANSWER

OpenStudy (amistre64):

24+n + 24 - n = 48 regardless of n ... good job

OpenStudy (anonymous):

@parik good alternate answer..:)